40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 15 22:16:09 2012, Jenne, Update, SUS, All oplev gains turned to 0 
    Reply  Mon Nov 19 15:15:22 2012, Jenne, Update, SUS, All oplev gains turned on 
Message ID: 7724     Entry time: Mon Nov 19 15:15:22 2012     In reply to: 7551
Author: Jenne 
Type: Update 
Category: SUS 
Subject: All oplev gains turned on 

Quote:

Oplev values that were changed to zero:

PRM P=0.15, Y=-0.3

SRM P=-2.0, Y=2.0

BS P=0.2, Y=-0.2

ITMY P=2.1, Y=-2.0

ITMX P=1.0, Y=-0.5

ETMX P=-0.2, Y=-0.2

ETMY P=0.5, Y=0.6

Also, PRCL was changed in the LSC input matrix from REFL33I to AS55I, since there is no REFL beam out of the IFO :(

 Ayaka and I restored all of the oplev gains to these values.  The exception is ETMY, which has both gains negative.  I am unsure if this is a transcription error on my part, or if something physical has changed.  The layout of the ETMY oplev was modified (since Rana took out the offending lens) but that shouldn't affect the sign of the gains.

ELOG V3.1.3-